OpenStudy (anonymous):

Please help me solve this! A tank is filled with 100L of a 40% alcohol solution (by volume) you repeatedly perform the following: remove 2 L of the soln from the tank and add 2L of 10 % alcohol soln A. let Cn be a concentration of the soln in the tank after the nth replacement where C0 =40% write the first five terms of {Cn} B. after how many replacements does the alcohol concentration reach 15%? thank you in advance!

Miracrown (miracrown):

Ok, well it seems like we have a recursive relation here. What have you been able to do with this problem so far?

OpenStudy (anonymous):

well I came up with Cn=40 - o.60n

OpenStudy (anonymous):

I don't think its correct

Miracrown (miracrown):

Sure, can you write out how you found that solution?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

c1=40L- 2L*40%+2L*10%= a1 c2= a1-2L*40%+2L*10% c3= a2-*40%+2L*10% cn=40L-n(2(40%-10%))

Miracrown (miracrown):

Ok, so first off the tank is 100L, not 40L and to simplify you could write 100L @40% - 2L @40% as 98L @ 40% then you add 2L at 10% correctly, but what is your new concentration going to be?

OpenStudy (anonymous):

but 100 L is 40% alcohol so i multpiplied

Miracrown (miracrown):

What you have there is *total* ethanol (in L), not concentration. so you want to divide by the total volume, yes?

OpenStudy (anonymous):

yes

Miracrown (miracrown):

Right, so 100L x 0.4 = 40L Ah, oh, hold on

Miracrown (miracrown):

Oh, it looks like you are figuring out total volume of ethanol... You can do it that way too. :D

OpenStudy (anonymous):

but i dont understand

Miracrown (miracrown):

Well, I suggest this: because it asks when the concentration drops below 15%, we should write a recursive relation for the concentration. just so that for each step we don't need to do an additional calculation to find the concentration. how does that sound?

OpenStudy (anonymous):

yah ok its a little challenging for me sorry

Miracrown (miracrown):

No worries. What I meant was, the expression you wrote would be correct if you were looking at just the total volume of ethanol. (how many of the 100L is ethanol). The way I suggested we do it is that we don't look at the *amount* but the *percentage of total* that is ethanol... because that is what the question is asking about. :)

OpenStudy (anonymous):

ok

Miracrown (miracrown):

Ok, so if we are to rewrite our recursive relationship in terms of %ethanol, we might end up with something like....

Miracrown (miracrown):